Question 1

Consider a single price monopolist that has standard U-shaped cost curves. If the monopolist earns positive economic profit, which of the following statements is true?

Group of answer choices

The monopolist sets a price greater than marginal cost.

The monopolist sets a price greater than average total cost.

The monopolist chooses level of output where marginal revenue is equal to marginal cost.

a, b and c are correct.

none of the above.

Question 2

Consider the following game. Terry and Kerry are roommates who must make decisions about cleaning simultaneously. Terry’s payoff is the first number in each cell and a higher number is a better outcome:

Kerry

Clean

Don’t clean

Terry

Clean

( 8, 2 )

( 3, 5 )

Don’t clean

( 10, 3 )

( 4, 1 )

Group of answer choices

This game has no Nash equilibrium.

This game has a single Nash equilibrium.

This game has more than one Nash equilibrium.

Both Terry and Kerry have a strictly dominant strategy.

both b and d are correct.

Question 3

Consider the following game. Terry and Kerry are roommates who must make decisions about cleaning simultaneously. Terry’s payoff is the first number in each cell and a higher number is a better outcome:

Kerry

Clean

Don’t clean

Terry

Clean

( 8, 2 )

( 3, 5 )

Don’t clean

( 10, 3 )

( 4, 1 )

Group of answer choices

Only Terry has a strictly dominant strategy.

Only Kerry has a strictly dominant strategy.

Both Terry and Kerry have a strictly dominant strategy.

Neither Terry and Kerry have a strictly dominant strategy.

None of the above.

Question 4

Consider the following game represented in normal form. Terry and Kerry are roommates who must make decisions about cleaning. Terry’s payoff is the first number in each cell and a higher number is a better outcome. Assume that Terry makes his decision first, if so what is the Nash Equilibrium of this game:

Kerry

Clean

Don’t clean

Terry

Clean

( 8, 2 )

( 3, 5 )

Don’t clean

( 10, 3 )

( 4, 1 )

Group of answer choices

Terry will clean and Kerry will not clean.

Terry will not clean and Kerry will not clean.

Terry will not clean and Kerry will clean.

Terry will clean and Kerry will clean.

More information is required to answer this question.

Question 5

A monopolist faces a demand curve given by Q=200-2p and has constant marginal (and average total cost) of 20. What is the economic profit made by this profit-maximising monopolist if they engage in perfect price discrimination?

Group of answer choices

0

800

3200

6400

None of the above

Question 6

A single price monopolist faces a demand curve given by Q=200-2p and has constant marginal (and average total cost) of 20. What is the economic profit made by this profit-maximising monopolist?

Group of answer choices

0

800

3200

6400

None of the above

Question 7

A single price monopolist faces a demand curve given by Q=200-2p and has constant marginal (and average total cost) of 20. What is the value of the deadweight loss generated by this monopolist?

Group of answer choices

0

800

3200

6400

None of the above

Question 8

Two shops in Newtown, “Malcolms Meatballs” and “Bill’s Burgers”, are deciding whether to hire a security guard. The security guard will provide deterrence to both stores. Each store will benefit from the presence of a security guard by reducing theft. The payoff matrix is given by and assume that Bill and Malcolm make their decision simultaneously and that a higher number represents a better outcome:

Bill’s Burgers
Hire Do not hire
Malcolm’s Meatballs Hire 6, 4 -2, 12
Do not hire 12, -6 -3, -3



In the Nash equilibrium:Group of answer choices

Bill hires a guard but Malcolm doesn’t.

Bill doesn’t hire a guard but Malcolm does.

Neither Bill nor Malcolm hire a guard.

Both Bill and Malcolm hire a guard.

More information is required to answer the question.

Question 9

Assume that a single price monopolist has a marginal cost curve given by MC=10+2Q. Further the demand curve that it faces is given by p=250-Q. Compared to a perfectly industry, the deadweight loss generated by this monopolist equals:

Group of answer choices

0.

600:

1800.

3600.

None of the above.

Question 10

Assume that individuals are homogeneous and that each has a demand curve of the following form for internet service: p=50-2q where p is the price per hour and q is hours per month. Assume the firm has a constant marginal cost of $12. The profit maximising two-part tariff results in the firm selling ______ hours and receiving total revenue of ________ from each consumer:

Group of answer choices

9.5: 361.

9.5: 589:

19: 361.

19: 589.

None of the above.
Answer #1

It's mandatory to answer only first four MCQ

Q1) option D)

All are right

Bcoz in monopoly, at eqm

MR = MC & price found from demand Curve

So P > MC

So a is true

For positive profit, P > ATC , b is true

2) option B)

Clean dont clean
Clean (8,2) (3,5•)
Dont clean (10*,3•) (4*,1)

NE : ( dont clean, clean )

Option B is right , only single NE exists

Only Row player Terry, dominant strategy : Don't clean

Kerry doesn't have the dominant strategy

3) option A)

Only Terry has a dominant strategy, bcoz only Terry gets higher payoff from don't clean always, for any strategy choice of Kerry

.

​​​​​​

4) option C)

Sequential game , so using game tree to solve the game .

Terry don't clean , Kerry will clean

5) option D) 6400

Perfect price discrimination, when P charged = MC

& Entire CS is taken as monopoly profits

So, Q = 200-2*20 = 160

P = MC = 20

So CS = .5*(100-20)*160 = 6400

.

6) option C)

Q = 200-2P

P = 100- Q/2

At eqm, MR = MC

100-Q = 20

Q* = 80

P = 100-40 = 60

π = (60-20)*80 = 3200